The first term of an Ap is 11 while the last term is 144. If the common difference is 7 find the number of term it has
Answer by mananth(16946)

a=11
d=7
tn = 144
tn=a+(n-1)d
144=11+(n-1)*7
144=11+7n-7
144-11+7=7n
140=7n
n= 140/7
n=20